# Towards Multi-Modal Animal Pose Estimation: An In-Depth Analysis

### v1

@@ 4.3.1. Single Camera View
Skeleton-only methods are dedicated to estimating the 3D coordinates of animal joints for a skeletal representation as their result. Inspired by the advancements in 2D APE, the paradigm of inferring 3D poses from estimated 2D poses, termed 2D-to-3D lifting, has garnered considerable attention. This approach typically leverages strong priors for predicting 3D joint locations from a single-camera view. A notable example is the LiftPose3D framework (Gosztolai et al., 2021), which trains a supervised neural network utilizing 3D ground truth data to lift 2D poses from one camera. The training data is augmented with different camera angles and bone lengths, which allows the network to internalize camera parameters implicitly and accommodate variations in animal sizes. In contrast, recent innovations have endeavoured to estimate camera parameters and construct shape distributions explicitly (Hu et al., 2023). In this paper (Hu et al., 2023), 3D APE is formulated as an optimization problem utilizing a pre-defined mouse skeletal model. This 3D skeleton model imposes a strong prior, addressing the issue of occlusions and acting as a regularization on the over-parameterized joint space. These two methods hinge on extensive 3D annotated datasets, which are hard to apply to wild animals whose data are difficult to collect.
@@
Recent attempts have been made to address this challenge to eliminate the dependence on 3D ground truth through unsupervised 2D-to-3D lifting techniques. For instance, Dai et al. (Dai et al., 2023) have introduced an unsupervised approach that, given a 2D pose, employs a pose generator network to generate a corresponding 3D posture, while a separate camera network estimates the camera rotation. During training, the generated 3D pose is subjected to random re-projections onto camera viewpoints to synthesize a new 2D pose. The synthesized 2D pose is then decomposed into a 3D pose and a camera rotation, based on which consistency losses are imposed in both 3D canonical poses and camera rotations for self-supervised training with geometric consistency. Complementing this, Sosa et al. (Sosa and Hogg, 2023) have proposed a similar methodology which relies on self-supervision with the same geometric consistency constraint for 3D pose estimation. Notably, while Dai et al. (Dai et al., 2023) assume the availability of actual 2D poses for each input image, Sosa et al. (Sosa and Hogg, 2023) demonstrate the feasibility of their approach with minimal assumptions, including unlabeled images and a small set of synthetic 2D poses. The minimalistic requirements and the availability of large volumes of unlabeled data suggest that this method (Sosa and Hogg, 2023) could be readily generalized across diverse animal species. Another 2024 approach to automatically extract the 3D pose from single-camera-view images using only synthetic data for training is DigiDogs (Shooter et al., 2024). This method takes DINOv2 (Oquab et al., 2023) as the base architecture, which is a foundation model known for its robust pre-training capabilities and superior domain adaptation performance. These studies exploit synthetic data to predict 3D joint locations from 2D images, reinforcing the potential of synthetic augmentation in the domain of monocular 3D APE.
@@
Animal Mesh Recovery
In the domain of 3D single-camera-view APE from monocular images and videos, this AMR subfield is further categorized into model-based AMR and model-free AMR. The former predominantly utilizes parametric body models to infer the detailed 3D mesh of animals. The latter eliminates the need for a pre-existing parametric template model to recover 3D animal meshes.
One of the most well-known model-based AMR is the Skinned Multi-Animal Linear (SMAL) model (Zuffi et al., 2017), which provides a template for various animal classes based on 3D scanning of deformable toy animals. The SMAL model represents animals using a linear combination of shape bases, which are adapted to capture the diversity of animal forms, alongside a set of pose parameters to articulate the model. This framework serves as a foundation for many model-based AMR works. A notable enhancement of the SMAL model is SMALR (Zuffi et al., 2018), which allows the recovery of detailed 3D shapes of unseen animals and creates textured 3D meshes. SMALST (Zuffi et al., 2019a) further amalgamates real textures with synthetically generated poses, shapes, and backgrounds. Beyond SMALST, the field has seen the development of breed-specific variants of SMAL, such as models for dogs (Rüegg et al., 2023, 2022; Biggs et al., 2020; Li and Lee, 2021a) and birds (Wang et al., 2021a). These models incorporate scaling parameters into the SMAL framework to enable breed-specific fitting. By using the pre-defined 3D shape and pose priors for the target animals, model-based AMR methods aid the animal shape reconstructions in ambiguous cases; the recovered mesh model further provides details for analyzing body size, posture, motion, behaviour, and health conditions. However, when we need to estimate the poses for new animal types without an existing template mesh, data collection for these unseen animals can be time-intensive, thus restricting model-based AMR’s applicability to the diverse animal kingdom.
On the other hand, model-free approaches (Kanazawa et al., 2018; Goel et al., 2020; Wu et al., 2023a, b; Yang et al., 2022a; Li et al., 2024b; Jakab et al., 2024; Yang et al., 2023b) do not rely on pre-existing parametric body models, which offers greater flexibility when we have rare and diverse animals. Many of these methods attempt to learn 3D priors from large-scale internet images and videos in an unsupervised or self-supervised manner, focusing on one category at a time (Goel et al., 2020; Wu et al., 2023a, b; Jakab et al., 2024; Yang et al., 2023b). In contrast, 3D-Fauna (2024) (Li et al., 2024b) introduces a pan-category deformable model capable of reconstructing any quadruped animal instance in a feed-forward manner, pushing towards a more generalizable and unconstrained understanding of animal forms from monocular images. However, the recovered 3D shapes of current model-free approaches often produce overly smooth surfaces, which do not preserve as much detail as the model-based methods; this severely limits their applicability for downstream applications.
4.3.2. Multiple Camera Views
@@
The majority of the existing methods in multiple-camera-views 3D APE are designed to estimate the pose of a single animal in keypoints format (Yao et al., 2019; Günel et al., 2019; Nath et al., 2019; Zhang and Park, 2020; Bala et al., 2020; Zhang et al., 2021; Dunn et al., 2021; Karashchuk et al., 2021; Muramatsu et al., 2022; Ebrahimi et al., 2023). In particular, early methods estimate the 2D poses from multiple independent camera views first; then, these 2D predictions are triangulated to 3D poses (LABEL:fig:multiview). These post hoc triangulation approaches like (Nath et al., 2019) work well when all the joints are visible in all views. However, they are not suited to handle occlusions, which is a scenario that always happens in animal naturalistic environments. To address these limitations, some methods apply skeletal constraints and filtering (Günel et al., 2019; Karashchuk et al., 2021). In detail, Anipose (Karashchuk et al., 2021) benefits significantly from post-processing spatial-temporal filtering, especially in scenarios involving freely moving subjects where occlusion is frequent. However, these methods have yet to achieve robust 3D tracking of freely moving animals unless trained with large numbers of frames. A notable exception is OpenMonkeyStudio (Bala et al., 2020), which can estimate the 3D macaques pose in extensive, unconstrained environments with a network of 62 high-resolution video cameras. However, these methods triangulate 3D locations from 2D estimates; their 2D nature limits their ability to capture and quantify behaviours in 3D accurately. To incorporate advances in 3D, some approaches integrate multiple-camera-views geometric data directly within end-to-end training pipelines (LABEL:fig:multiview2). A notable system, DANNCE (Dunn et al., 2021), utilizes projective geometry to construct a 3D feature space from multiple-camera-views imagery and subsequently trains a 3D CNN to predict the positions of anatomical landmarks from the shared features across cameras. In addition, to tackle the data scarcity in 3D APE, 3D-Upper (2023) (Ebrahimi et al., 2023) uses an unsupervised algorithm based on a Statistical Shape Model (SSM) to estimate 3D poses of freely moving animals; this SSM estimator is robust even when many poses (up to 50%) are contaminated by outliers and missing data. The 3D-Upper APE method does not require labelled training data, which makes it highly scalable and more practical for applications in-the-wild or in situations where manual annotations are impractical.
@@
All of the aforementioned 3D multiple-camera-views approaches are limited to estimating the pose of a single animal, and they only provide poses in keypoints format, which does not preserve as much animal shape information as mesh-based APE methods. MAMMAL (2024) (An et al., 2023) and SBeA (2024) (Han et al., 2024) extend this to enable multi-animal 3D social pose estimation using multi-camera-views. In particular, MAMMAL is designed for large mammals and provides a full 3D surface mesh model compared to the sparse 2D keypoint methods. It captures fine-level surface geometries to understand animal motion involving interactions and occlusions in their complex social behaviours. Moreover, the cross-view matching algorithm further allows MAMMAL to reconstruct 3D poses robustly, even in the presence of keypoint occlusion. Triangulation-based methods are weak in such scenarios and may lead to incomplete or incorrect pose estimations. Another recent method, SBeA, further pushes the trend towards using a few-shot learning strategy to accurately estimate multi-animal 3D poses and identities with minimal manual annotation. SBeA introduces a novel bidirectional transfer learning strategy that enables zero-shot multi-animal identity recognition. It reuses models trained on single-animal data to identify multiple animals, thus minimizing extra identity annotations. The SBeA model employs multiview camera arrays and geometrical constraints to reconstruct 3D multiple animals’ poses effectively, which is hard to attain through single-camera-view APE methods. Lastly, SBeA confers the capability for unsupervised behaviour classification by decomposing 3D trajectories into 2D representations and clustering social behaviour modules. This permits the discovery of subtle social behaviour phenotypes that supervised techniques might miss.
@@
In summary, RGB-based APE methods have advanced significantly from 2017 to 2024. Early approaches, like DeepLabCut (Mathis et al., 2018), focused on 2D keypoint-based supervised methods adapted from HPE for tracking animal joints in controlled environments. As research progressed, interest grew in 3D APE (particularly mesh-based methods), which capture more detailed pose and shape data for downstream tasks like tracking and behaviour analysis. However, the limited annotated data, and variations in animal species and environmental set-up in APE have initially restricted the progress of supervised APE methods. To address these, recent methods have shifted to weakly-supervised, self-supervised, semi-supervised, and unsupervised learning, using synthetic data, transfer learning, and domain adaptation to improve model generalization and tackle data scarcity. Additionally, given that many animals are social creatures, there is a trend toward multi-animal tracking from video data to understand their communication and social structures. These trends culminated in 2024 with the release of SuperAnimal (Shaokai Ye and Mathis, 2024) (2D APE), MAMMAL (An et al., 2023) and SBeA (Han et al., 2024) (3D APE), which offer superior APE performance, generalization across species and diverse environments, reduced dependence on labelled data, and allow for tracking multiple animals simultaneously, setting a new benchmark in the field.
5. Non-RGB-based Unimodal Animal Pose Estimation
@@
In the previous section, we comprehensively reviewed APE methods primarily based on RGB cameras. However, beyond RGB-based approaches, researchers have explored the use of other unimodal methods for analyzing animal behaviour, table 8 summarises the latest APE methods using both non-RGB unimodal (Yigit et al., 2022) and multi-modal approaches, which is discussed in detailed in section 6.3.
IMU
@@
IMU is a small, lightweight, low-cost device that measures an object’s acceleration and angular velocity. Attaching multiple IMUs to limbs and the trunk makes it possible to capture the motion and orientation of the animal in 3D space. Recent work (Yigit et al., 2022) presents a lameness detection and pose estimation method for horse walk and trot locomotion based on the acceleration and gyroscope measurements from IMUs. Initially, they train the network offline using IMU measurements as input sequences, with labelled data provided by the MoCap system. Once the long short-term memory (LSTM) model for gait event and lameness detection is trained, it embeds the IMU measurements into high-dimensional joint angles, which are then fed into the Gaussian Process Dynamic Model (GPDM) to learn a low-dimensional latent representation of the horse’s limb motion. This learned latent manifold dynamics is parameterized with the gait phase variable to predict the horse’s 24 limb joints in real time.
@@
Table 8. Non-RGB-based unimodal APE methods and multi-modal APE methods. All the symbols and column meanings are defined in table 5, table 6 and table 7, and any additional symbols are defined here: [Uncaptioned image] is GPDM architecture, [Uncaptioned image] is Hierarchical Gaussian Process Latent Variable Model (H-GPLVM), [Uncaptioned image]is Extended Kalman Smoother (EKS), [Uncaptioned image] is Context Models, [Uncaptioned image] is CLIP. [Uncaptioned image] is ego-view. Ther. is an abbreviation for thermal, and Lang. is an abbreviation for language.
@@
1 2022 Yigit et al. (Yigit et al., 2022) IMU [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Horse Mean, Std, VR, IAE & IEV
@@
2 2017 Patel et al. (Patel et al., 2017) RGB+GPS+ [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Cheetah Qualitative for 3D
IMU (video observations)
3 2020 Kearney et al. (Kearney et al., 2020a) RGBD [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] RGBD-Dog (Kearney et al., 2020a), PCK, MPJPE, PA-PCK,
BADJA (Biggs et al., 2018) PCK 3D, PA-PCK 3D,
PA-MPJPE
4 2023 Luo et al. (Luo et al., 2023) RGBD {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Cattles, Pigs (Ruchay et al., 2020) MAE, MRE (Luo et al., 2023)
5 2023 OptiPose (Patel et al., 2023) RGBD+Ther. [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Rodent3D (Patel et al., 2023), PCK, MPJPE,
+Infrared Rat7M (Dunn et al., 2021), AcinoSet (Joska et al., 2021) Reprojection Error
@@
6 2024 WildPose(Muramatsu et al., 2024) RGB+LiDAR [Uncaptioned image], N/A, [Uncaptioned image], [Uncaptioned image] Eagle, Giraffe, Lion Mean & Std
7 2022 Sound of Motion (Li et al., 2022) RGB+Acoustic [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse PCK, IoU, PA-MPJPE
8 2024 CLHOP (Li et al., 2024a) RGB+Acoustic [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], {[Uncaptioned image],[Uncaptioned image]} Horse (Li et al., 2024a) PCK@0.1, IoU, PA-MPJPE
9 2023 CLAMP (Zhang et al., 2023) RGB+Lang. [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b), AP, AR
Animal Pose (Cao et al., 2019)
10 2023 AmadeusGPT (Ye et al., 2023) RGB+Lang. [Uncaptioned image], [Uncaptioned image], N/A, [Uncaptioned image] Mouse & others (Ye et al., 2023) F1 Score
11 2024 AWOL (Zuffi and Black, 2024) RGB+Lang. [Uncaptioned image], [Uncaptioned image], N/A, [Uncaptioned image] StanfordExtra (Biggs et al., 2020), Perceptual Study,
3D Trees CLIP score, Qualitative
6. Multi-Modal Animal Pose Estimation

### v2

@@ 4.3.1 Single Camera View
Skeleton-only methods are dedicated to estimating the 3D coordinates of animal joints for a skeletal representation as their result. Inspired by the advancements in 2D APE, the paradigm of inferring 3D poses from estimated 2D poses, termed 2D-to-3D lifting, has garnered considerable attention. This approach typically leverages strong priors for predicting 3D joint locations from a single-camera view. A notable example is the LiftPose3D framework [39], which trains a supervised neural network utilizing 3D ground truth data to lift 2D poses from one camera. The training data is augmented with different camera angles and bone lengths, which allows the network to internalize camera parameters implicitly and accommodate variations in animal sizes. In contrast, recent innovations have endeavoured to estimate camera parameters and construct shape distributions explicitly [47]. In this paper [47], 3D APE is formulated as an optimization problem utilizing a pre-defined mouse skeletal model. This 3D skeleton model imposes a strong prior, addressing the issue of occlusions and acting as a regularization on the over-parameterized joint space. These two methods hinge on extensive 3D annotated datasets, which are hard to apply to wild animals whose data are difficult to collect.
@@
Recent attempts have been made to address this challenge to eliminate the dependence on 3D ground truth through unsupervised 2D-to-3D lifting techniques. For instance, Dai et al. [19] have introduced an unsupervised approach that, given a 2D pose, employs a pose generator network to generate a corresponding 3D posture, while a separate camera network estimates the camera rotation. During training, the generated 3D pose is subjected to random re-projections onto camera viewpoints to synthesize a new 2D pose. The synthesized 2D pose is then decomposed into a 3D pose and a camera rotation, based on which consistency losses are imposed in both 3D canonical poses and camera rotations for self-supervised training with geometric consistency. Complementing this, Sosa et al. [132] have proposed a similar methodology which relies on self-supervision with the same geometric consistency constraint for 3D pose estimation. Notably, while Dai et al. [19] assume the availability of actual 2D poses for each input image, Sosa et al. [132] demonstrate the feasibility of their approach with minimal assumptions, including unlabeled images and a small set of synthetic 2D poses. The minimalistic requirements and the availability of large volumes of unlabeled data suggest that this method [132] could be readily generalized across diverse animal species. Another 2024 approach to automatically extract the 3D pose from single-camera-view images using only synthetic data for training is DigiDogs [128]. This method takes DINOv2 [105] as the base architecture, which is a foundation model known for its robust pre-training capabilities and superior domain adaptation performance. These studies exploit synthetic data to predict 3D joint locations from 2D images, reinforcing the potential of synthetic augmentation in the domain of monocular 3D APE.
@@
On the other hand, AMR for 3D APE from monocular images and videos is further categorized into model-based AMR and model-free AMR. The former predominantly utilizes parametric body models to infer the detailed 3D mesh of animals. The latter eliminates the need for a pre-existing parametric template model to recover 3D animal meshes.
One of the most well-known model-based AMR is the Skinned Multi-Animal Linear (SMAL) model [178], which provides a template for various animal classes based on 3D scanning of deformable toy animals. The SMAL model represents animals using a linear combination of shape bases, which are adapted to capture the diversity of animal forms, alongside a set of pose parameters to articulate the model. This framework serves as a foundation for many model-based AMR works. A notable enhancement of the SMAL model is SMALR [177], which allows the recovery of detailed 3D shapes of unseen animals and creates textured 3D meshes. SMALST [176] further amalgamates real textures with synthetically generated poses, shapes, and backgrounds. Beyond SMALST, the field has seen the development of breed-specific variants of SMAL, such as models for dogs [120, 121, 10, 68] and birds [144]. These models incorporate scaling parameters into the SMAL framework to enable breed-specific fitting. By using the pre-defined 3D shape and pose priors for the target animals, model-based AMR methods aid the animal shape reconstructions in ambiguous cases; the recovered mesh model further provides details for analyzing body size, posture, motion, behaviour, and health conditions. However, when we need to estimate the poses for new animal types without an existing template mesh, data collection for these unseen animals can be time-intensive, thus restricting model-based AMR’s applicability to the diverse animal kingdom.
On the other hand, model-free AMR approaches [54, 36, 153, 147, 154, 148, 74, 50] do not rely on pre-existing parametric body models, which offers greater flexibility when we have rare and diverse animals. Many of these methods attempt to learn 3D priors from large-scale internet images and videos in an unsupervised or self-supervised manner, focusing on one category at a time [36, 147, 154, 148, 50]. In contrast, 3D-Fauna (2024) [74] introduces a pan-category deformable model capable of reconstructing any quadruped animal instance in a feed-forward manner, pushing towards a more generalizable and unconstrained understanding of animal forms from monocular images. However, the recovered 3D shapes of current model-free approaches often produce overly smooth surfaces, which do not preserve as much detail as the model-based methods; this severely limits their applicability for downstream applications.
4.3.2 Multiple Camera Views
@@
(a) An example of multiple camera views 3D APE method [30] that encompasses multi-views approach based on triangulation of 2D keypoint estimations.
(b) An example of multiple camera views 3D APE method [30] that leverages multi-views geometric information during end-to-end training.
Figure 5: Multiple camera views 3D animal pose estimation using “DAN-NCE” [30].
@@
The majority of the existing methods in multiple-camera-views 3D APE are designed to estimate the pose of a single animal in keypoints format [159, 41, 100, 170, 8, 168, 30, 55, 96, 31]. In particular, early methods estimate the 2D poses from multiple independent camera views first; then, these 2D predictions are triangulated to 3D poses (Figure 5(a)). These post hoc triangulation approaches like [100] work well when all the joints are visible in all views. However, they are not suited to handle occlusions, which is a scenario that always happens in animal naturalistic environments. To address these limitations, some methods apply skeletal constraints and filtering [41, 55]. In detail, Anipose [55] benefits significantly from post-processing spatial-temporal filtering, especially in scenarios involving freely moving subjects where occlusion is frequent. However, these methods have yet to achieve robust 3D tracking of freely moving animals unless trained with large numbers of frames. A notable exception is OpenMonkeyStudio [8], which can estimate the 3D macaques pose in extensive, unconstrained environments with a network of 62 high-resolution video cameras. However, these methods triangulate 3D locations from 2D estimates; their 2D nature limits their ability to capture and quantify behaviours in 3D accurately. To incorporate advances in 3D, some approaches integrate multiple-camera-views geometric data directly within end-to-end training pipelines (Figure 5(b)). A notable system, DANNCE [30], utilizes projective geometry to construct a 3D feature space from multiple-camera-views imagery and subsequently trains a 3D CNN to predict the positions of anatomical landmarks from the shared features across cameras. In addition, to tackle the data scarcity in 3D APE, 3D-Upper (2023) [31] uses an unsupervised algorithm based on a Statistical Shape Model (SSM) to estimate 3D poses of freely moving animals; this SSM estimator is robust even when many poses (up to 50%) are contaminated by outliers and missing data. The 3D-Upper APE method does not require labelled training data, which makes it highly scalable and more practical for applications in-the-wild or in situations where manual annotations are impractical.
@@
All of the aforementioned 3D multiple-camera-views approaches are limited to estimating the pose of a single animal, and they only provide poses in keypoints format, which does not preserve as much animal shape information as mesh-based APE methods. MAMMAL [2] and SBeA [43] extend this to enable multi-animal 3D social pose estimation using multi-camera-views. In particular, MAMMAL is designed for large mammals and provides a full 3D surface mesh model compared to the sparse 2D keypoint methods. It captures fine-level surface geometries to understand animal motion involving interactions and occlusions in their complex social behaviours. Moreover, the cross-view matching algorithm further allows MAMMAL to reconstruct 3D poses robustly, even in the presence of keypoint occlusion. Triangulation-based methods are weak in such scenarios and may lead to incomplete or incorrect pose estimations. Another recent method, SBeA, further pushes the trend towards using a few-shot learning strategy to accurately estimate multi-animal 3D poses and identities with minimal manual annotation. SBeA introduces a novel bidirectional transfer learning strategy that enables zero-shot multi-animal identity recognition. It reuses models trained on single-animal data to identify multiple animals, thus minimizing extra identity annotations. The SBeA model employs multiview camera arrays and geometrical constraints to reconstruct 3D multiple animals’ poses effectively, which is hard to attain through single-camera-view APE methods. Lastly, SBeA confers the capability for unsupervised behaviour classification by decomposing 3D trajectories into 2D representations and clustering social behaviour modules. This permits the discovery of subtle social behaviour phenotypes that supervised techniques might miss.
@@
In summary, RGB-based APE methods have advanced significantly from 2011 to 2024. Early approaches, like DeepLabCut [90], focused on 2D keypoint-based supervised methods adapted from HPE for tracking animal joints in controlled environments. As research progressed, interest grew in 3D APE (particularly mesh-based methods), which capture more detailed pose and shape data for downstream tasks like tracking and behaviour analysis. However, the limited annotated data, and variations in animal species and environmental set-up in APE have initially restricted the progress of supervised APE methods. To address these, recent methods have shifted to weakly-supervised, self-supervised, semi-supervised, and unsupervised learning, using synthetic data, transfer learning, and domain adaptation to improve model generalization and tackle data scarcity. Additionally, given that many animals are social creatures, there is a trend toward multi-animal tracking from video data to understand their communication and social structures. These trends culminated in 2024 with the release of SuperAnimal [160] (2D APE), MAMMAL [2] and SBeA [43] (3D APE), which offer superior APE performance, generalization across species and diverse environments, reduced dependence on labelled data, and allow for tracking multiple animals simultaneously, setting a new benchmark in the field.
5 Non-RGB-based Unimodal Animal Pose Estimation
@@
In the previous section, we comprehensively reviewed APE methods primarily based on RGB cameras. However, beyond RGB-based approaches, researchers have explored the use of other unimodal methods for analyzing animal behaviour, Table 8 summarises the latest APE methods using both non-RGB unimodal [162] and multi-modal approaches, which is discussed in detailed in Section 6.3.
@@
IMU: IMU is a small, lightweight, low-cost device that measures an object’s acceleration and angular velocity. Attaching multiple IMUs to limbs and the trunk makes it possible to capture the motion and orientation of the animal in 3D space. Recent work [162] presents a lameness detection and pose estimation method for horse walk and trot locomotion based on the acceleration and gyroscope measurements from IMUs. Initially, they train the network offline using IMU measurements as input sequences, with labelled data provided by the MoCap system. Once the long short-term memory (LSTM) model for gait event and lameness detection is trained, it embeds the IMU measurements into high-dimensional joint angles, which are then fed into the Gaussian Process Dynamic Model (GPDM) to learn a low-dimensional latent representation of the horse’s limb motion. This learned latent manifold dynamics is parameterized with the gait phase variable to predict the horse’s 24 limb joints in real time.
@@
Table 8: Non-RGB-based unimodal APE methods and multi-modal APE methods. All the symbols and column meanings are defined in Table 5, Table 6 and Table 7, and any additional symbols are defined here: [Uncaptioned image] is GPDM architecture, [Uncaptioned image] is Hierarchical Gaussian Process Latent Variable Model (H-GPLVM), [Uncaptioned image]is Extended Kalman Smoother (EKS), [Uncaptioned image] is Context Models, [Uncaptioned image] is CLIP. [Uncaptioned image] is ego-view. Ther. is an abbreviation for thermal, and Lang. is an abbreviation for language. The method citation is available in column Year.
@@
1 2022 Yigit et al. IMU [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Horse Mean, Std, VR, IAE & IEV
@@
2 2017 Patel et al. RGB+GPS+IMU [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Cheetah Qualitative for 3D
(video observations)
3 2020 Kearney et al. RGBD [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] BADJA (2018), PCK, MPJPE, PA-PCK,
RGBD-Dog (2020) PA-MPJPE, PCK 3D,
PA-PCK 3D
4 2023 Luo et al. RGBD {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Cattles, Pigs (2020) MAE, MRE (2023)
5 2023 OptiPose RGBD+Ther. [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Rat7M (2021), PCK, MPJPE,
Infrared AcinoSet (2021), Reprojection Error
Infrared Rodent3D (2023)
@@
6 2024 WildPose RGB+LiDAR [Uncaptioned image], N/A, [Uncaptioned image], [Uncaptioned image] Eagle, Giraffe, Lion Mean & Std
7 2022 Sound of Motion RGB+Acoustic [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse PCK, IoU, PA-MPJPE
8 2024 CLHOP RGB+Acoustic [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], {[Uncaptioned image],[Uncaptioned image]} Horse (2024) PCK@0.1, IoU, PA-MPJPE
9 2023 CLAMP RGB+Lang. [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Animal Pose (2019), AP, AR
AP-10K (2021b)
10 2023 AmadeusGPT RGB+Lang. [Uncaptioned image], [Uncaptioned image], N/A, [Uncaptioned image] Mouse & others (2023) F1 Score
11 2024 AWOL RGB+Lang. [Uncaptioned image], [Uncaptioned image], N/A, [Uncaptioned image] StanfordExtra (2020), CLIP score, Qualitative,
3D Trees Perceptual Study
6 Multi-Modal Animal Pose Estimation